OpenStudy (anonymous):

Why does lim (1 + 1/x)^x when x tends to + inf, equals e?

OpenStudy (anonymous):

You know \[e ^{\ln x}\] is always equal to x. Just try doig the same thing there and you'll see what will happen.

OpenStudy (anonymous):

I really don't get it, would you mind to explain a bit more? Should I do: \[\lim_{x \rightarrow \infty} e^{\ln (1+1/x)^x} \]

OpenStudy (anonymous):

Perhaps by definition. Let that limit equal a value we call e.

OpenStudy (zarkon):

you would still need to prove that the limit converges.